A boy on a bridge throws a stone horizontally with a speed of 25 m/s releasing the stone from a point 19.6 m above the surface of the river. How far from a point directly below the boy will the stone strikes the water?


1
Expert's answer
2021-09-17T15:33:24-0400

Vertical and horizontal motions are independent. Since there is no forces acting on the stone in horizontal direction, its velocity in this direction does not change "v_h = v_{0h} = 25m\/s".

Time of its fall in vertical direction can be found from the kinematic equation:


"h =\dfrac{gt^2}{2}"

where "h = 19.6m" is the height of the surface, "g = 9.81m\/s^2" is the gravitational acceleration. Thus, obtian:


"t = \sqrt{\dfrac{2h}{g}}"

In this time the stone covers the following horizontal distance:


"d = tv_h =v_h \sqrt{\dfrac{2h}{g}}\\
d = 25m/s \sqrt{\dfrac{2\cdot 19.6m}{9.81m/s^2}} \approx 50m"

Answer. 50 m.